OpenStudy (anonymous):

Which of the following is the formula for this geometric sequence? 1, -3, 9, -27,... A.an = -1(3)n-1 B.an = 1(-3)n-1 C.an = 3(-1)n-1

Parth (parthkohli):

an = 1(-3)n-1. Can I explain or is that enough?

OpenStudy (anonymous):

explain is a must to support your answer...

Parth (parthkohli):

Okay. (-3)^1 = -3 (n - 1 = 1 as n = 2) This is an example.